OpenStudy (anonymous):

square root -100 x square root -81

OpenStudy (anonymous):

\[\sqrt{100}=10\] \[\sqrt{81}=9\] 10*9=90

OpenStudy (anonymous):

but doesn't the negative sign affect the answer?

OpenStudy (anonymous):

well then there wasn't an answer because you can't square root negatives

OpenStudy (anonymous):

I don't think you can sqrt negative even numbers, unless you are going into the imaginary number realm? i?

OpenStudy (anonymous):

yeah well would it be 90i ?

jimthompson5910 (jim_thompson5910):

\[\Large \sqrt{-100}\times\sqrt{-81}\] \[\Large \sqrt{-1*100}\times\sqrt{-1*81}\] \[\Large \sqrt{-1}\times\sqrt{100}\times\sqrt{-1}\times\sqrt{81}\] \[\Large i\times10\times i\times9\] \[\Large 90i^2\] \[\Large 90(-1)\] \[\Large -90\] So \[\Large \sqrt{-100}\times\sqrt{-81}=-90\]

OpenStudy (anonymous):

Thank you!